Resolve the following vectors into their components:

a)7.5m/s2 [NW] b)15m/s[25*WofS]


1
Expert's answer
2021-02-05T02:13:55-0500

a)

"a_x=7.5cos\theta, a_y=7.5sin\theta."

b)

"v_x=15cos(180^{\circ}+25^{\circ})=-13.59\ \dfrac{m}{s},""v_y=15sin(180^{\circ}+25^{\circ})=-6.34\ \dfrac{m}{s}."